Two cars leaves denver at the same time and travel in opposite directions. one car travels 10 mi/h faster than the other car. The cars are 500 mi apart in 5 h. How fast is each car traveling?
Answer by mananth(16946)

car speed = x
other car speed = x+10
..
The combined speed = x+x+10= 2x+10
D= 500 miles
rate = 2x+10
time = 5
500/(2x+10)=5
cross multiply
500 = 5(2x+10)
500=10x+50
10x=450
/10
x=45 mph
other car = x+10 = 55 mph